OpenStudy (anonymous):

Use the double angle formula to rewrite the equation y = 4sin^2 (slashed zero) - 2.

OpenStudy (anonymous):

is the slashed zero theta?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

it reduces to \[y = -2 \cos(2\theta)\]

OpenStudy (anonymous):

Can you show me how it works?

OpenStudy (anonymous):

You need to simply use that cos(2A) = 1- 2 sin^2(A)

OpenStudy (anonymous):

so from this we take theta to be A

OpenStudy (anonymous):

if you rewrite the equation you gave in that form it becomes.. -2 + 4sin^2(A)

OpenStudy (anonymous):

from here we can factor out a -2 which leaves us with -2(1-2sin^2(A)) and you can see that the values under the parenthesis is just cos(2A) from the identity. So the answer you end up with after plugging in theta for A, is just -2cos(2Theta)

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

Yes. Thank you for giving a detailed explanation.

OpenStudy (anonymous):

yeah no problem, glad i could help